# Robust Denoising Neural Reranker for Recommender Systems

### v1

@@ 3.3 Implementing Noise Injection Posterior
  • Heuristic Sampler: This option uses a non-parametric sampler $\bm{\epsilon}_{\phi}^{\text{heuristic}}$ with a manually designed distribution. For example, we can adopt standard Gaussian noises similar to [31, 26], but we notice that this is not a good representation of the ground truth. Intuitively, given the binary user response signals, a natural assumption of the retriever prior is the Beta distribution, where the parameters depend on the user request $\mathbf{u}$. What follows is a Beta posterior distribution, which derives our heuristic solution as $\bm{\epsilon}_{\phi}\sim\text{Beta}(\alpha,\beta)$. For simplicity, we set $\alpha=\beta=0.5$ to approximate the uncertainty, which would help improve the chance of generating harder samples for the reranker and consequently improve the robustness of the solution. Note that this implementation assumes that the noise does not depend on the user, and the personalization only comes from the label $\mathbf{z}_{u}$;
@@
  • Model-based Sampler: While the heuristic sampler is effective in our empirical study, we are skeptical on whether the fixed setting of $\alpha$ and $\beta$ is optimal for all user requests. To achieve a personalized noise estimation, we use a learnable model $\bm{\epsilon}_{\phi}^{\text{model}}$ with parameter $\phi$ that will be optimized along with the reranker (details in the next section). As we will show in Section 4.3, this choice is empirically better, and the learned error exhibits a Beta distribution without explicitly defining it. Without specification, our standard DNR solution uses this model-based sampler.
@@ 3.4 Overall Learning Framework
In this section, we show how to optimize the user feedback alignment i.e., Eq.(1) by collaborating the learning of the reranker and the discovery of better noise-injecting posterior. Specifically, following a variational inference framework, we can decompose the negative log-likelihood objective into three loss terms (detailed derivation in Appendix A.1):
$-\log p(\mathbf{z}_{u})$ $=\mathcal{L}_{\text{z}}+\mathcal{L}_{\text{x}}+\mathcal{L}_{\text{adv}}+\delta_{\text{x}}$ (5)
@@
which consists of an augmented feedback reconstruction loss term $\mathcal{L}_{\text{z}}$, an adversarial noise exploration loss $\mathcal{L}_{\text{adv}}$, and a posterior distribution regularization term $\mathcal{L}_{\text{x}}$. The additional term $\delta_{\text{x}}=-D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})\big)\leq 0$ represents the divergence between the noise injection posterior $p_{\phi}$ and the corresponding ground truth $p_{\text{x|z}}$. Intuitively, we expect to achieve a small divergence in $\delta_{\text{x}}$ and $L_{\text{x}}$ so that the optimization of the three terms is aligned with the goal of Eq.(1).
Denoising with Augmented Retriever Scores: $\mathcal{L}_{\text{z}}$ in Eq.(5) denotes the reconstruction loss under the noisy retriever scores sampled by the posterior. Specifically, we can first sample $\mathbf{x}_{u}^{\prime}\sim p_{\phi}(\cdot|\mathbf{z}_{u})$ as the generated retriever scores, then pass them into the reranker $q_{\theta}$ for the reconstruction of $\mathbf{z}_{u}$. The optimization can also be simplified into the BCE loss with $\mathbf{z}_{u}$ as labels, and it only updates $\theta$. Note that the generated scores $\mathbf{x}_{u}^{\prime}$, different from the observed $\mathbf{x}_{u}$, essentially serve as the augmented scores that the initial retriever would have produced given the user request. This means that the reranker will have to learn a more robust reconstruction model of the ground-truth user feedback. Combining with the direct optimization, we have the final reconstruction loss for the reranker:
@@
where $\lambda_{m}$ controls the magnitude of the augmentation, and the posterior $p_{\phi}$ is fixed.
Adversarial Noise Learning: $\mathcal{L}_{\text{adv}}$ describes how the reranker behaves differently from the true user feedback under the viewpoint of the noise-injecting posterior $p_{\phi}$. In practice, this term is expected to be small since the ratio $q_{\theta}(\mathbf{z}_{u}|\mathbf{x})/p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})\approx 1$ in major cases due to the user feedback alignment of Eq.(6) and the augmented objective $\mathcal{L}_{\text{z}}$. Nevertheless, in the view of the posterior, we can see that $\mathcal{L}_{\text{adv}}$ encourages adversarial samples of $\mathbf{x}$ that produce small ratios of $q_{\theta}(\mathbf{z}_{u}|\mathbf{x})/p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})$. In other words, a better $p_{\phi}$ should be able to find irregular samples that are correctly recovered by the true distribution $p_{\text{z|x}}$ but are incorrectly reconstructed by the reranker $q_{\theta}$. For the posterior with heuristic sampler $\bm{\epsilon}_{\phi}^{\text{heuristic}}$, it is not optimizable once $\alpha$ and $\beta$ are chosen. In contrast, for the model-based sampler $\bm{\epsilon}_{\phi}^{\text{model}}$, we follow the intuition of irregular case exploration, and approximate $\mathcal{L}_{\text{adv}}$ by first fixing the reranker and then optimizing the posterior with a reconstruction minimization loss:
@@
In our design, we directly output $\mathbf{x}_{u}^{\prime}$ using $\phi$ so that the end-to-end learning becomes feasible for Eq.(7). In cases where the sampling process is not directly optimizable, one may have to consider other alternatives like reparameterization tricks [32] or policy gradient [33].
On the other hand, $\mathcal{L}_{\text{adv}}$ does not encourage over-exploration of $\mathbf{x}_{u}^{\prime}$, which may negatively impact the learning of the reranker. In such cases, the generated scores are unrealistically different from the true posterior, diverging from the user behavior, inducing small ground truth $p_{\text{z|x}}$ and large ratio of $q_{\theta}(\mathbf{z}_{u}|\mathbf{x})/p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})$, violating $\mathcal{L}_{\text{adv}}$.
Noise Regularization: $\mathcal{L}_{\text{x}}$ estimates the distributional divergence between the posterior estimator $p_{\phi}$ and the prior $p_{\text{x}}$, and we aim to find a $p_{\phi}$ that reduces this divergence. Intuitively, this term favors the posteriors that can mimic the noisy behavior of the prior $p_{\text{x}}$ and remove the additional accuracy-related information from $\mathbf{z}$. We can calculate this term by first sampling retriever scores as $\mathbf{x}^{\prime}_{u}\sim p_{\phi}(\cdot|\mathbf{z}_{u})$, then estimating their log ratio difference from that of the prior. In cases where we do not have the model to directly estimate $p_{\text{x}}(\mathbf{x}_{u}^{\prime})$, one can circumvent this issue by directly aligning $\mathbf{x}_{u}^{\prime}$ with the observed samples $\mathbf{x}_{u}$:
$\mathcal{L}_{\text{x}}^{\prime}\approx-\mathbb{E}_{p_{\phi}}\big[\log\frac{\mathbf{x}_{u}}{\mathbf{x}^{\prime}_{u}}\big],$ (8)
which is an instance-level objective more rigorous than $\mathcal{L}_{\text{x}}$. Intuitively, a smaller value of $\mathcal{L}_{\text{x}}^{\prime}$ means that $\phi$’s output is close to the prior’s output $\mathbf{x}_{u}$, which restricts the distribution-level objective $\mathcal{L}_{\text{x}}$ to small values. Yet, we remind readers that it is not the case vice versa. Additionally, it is infeasible to repeat the exact same request and obtain augmented samples to cover the prior distribution of $p_{\text{x}}$. Fortunately, the collaborative learning across different user requests will approximate the sampling process of pair $(\mathbf{x}^{\prime}_{u},\mathbf{x}_{u})$ with similar patterns.
Again, note that the heuristic sampler $\bm{\epsilon}_{\phi}^{\text{heuristic}}$ is not optimizable by this objective, but one can achieve a small value of $\mathcal{L}_{\text{x}}$ by carefully designing the distribution (e.g., choosing Beta instead of Gaussian). For the training of the posterior with model-based sampler $\bm{\epsilon}_{\phi}^{\text{model}}$, we first adopt the heuristic sampler in the first $\lambda_{e}$ epochs to ensure learning stability, then engage the adversarial learning of $\mathcal{L}_{\text{adv}}$ and the optimization of Eq.(8) afterwards, switching the sampler to $\bm{\epsilon}_{\phi}^{\text{model}}$. As the summary of our solution, we present our training procedure in Appendix B.
@@ 4 Experiments
RQ1: How does DNR perform in the reranking tasks of recommendation systems compared with leading baselines? RQ2: How well is the denoising formulation compared to alternative methods of leveraging retriever scores? RQ3: What is the performance difference between Gaussian noise and Beta noise? RQ4: How does the model-based sampler compare to the heuristic sampler in practice? RQ5: How does different objectives in DNR contribute to the performance? RQ6: How sensitive is DNR to the hyperparameters $\lambda_{c}$, $\lambda_{m}$, and $\lambda_{e}$?
@@ 4.1 Expermental Settings
Datasets. We conduct experiments on three real-world datasets, including ML-1M [34], Kuaivideo [3], and Amazon-books [35]. The detailed description and the statistics of the processed datasets for both the retriever and the reranker stages are presented in Appendix C.1.
Baselines: We compare the performance of DNR with four categories of recommendation methods in our reranking task, including traditional recommenders (i.e., SASRec [23], Caser [36], GRU4Rec [37], and MiDNN [24]), which individually predict the scores of candidate items and rank them accordingly; list-refinement methods (i.e., DLCM [8], SetRank [38], PRM [6], and MIR [25]), which encode the mutual inference in candidate items and refine the rankings based on this information; generator-evaluator methods (i.e., EGRerank [22], Pier [9], and NAR4Rec [10]), which generate multiple reranked item lists and evaluate them to select the best one for users; and diffusion-Based methods (i.e., DiffuRec [26] and DCDR [11]) which utilize diffusion models to generate items or lists for recommendation. Detailed information for our baselines is presented in Appendix C.2.
Implementation Details: Our experiments are implemented on Nvidia A40 GPUs with Python 3.8 and PyTorch 1.12. We construct the retriever stage with the collaborative filtering method [39] to obtain the initial retriever scores for the top-50 candidate items. The detailed setting and experimental results of the retriever stage are presented in Appendix D.1. The reranker reorders these items filtered by retrievers and selects the top $K=6$ for display. The dimension of rerankers’ hidden embedding is 128 across all models, and all models are trained until convergence. The learning rate and weight decay are tuned within the range of $[0.01,0.001,0.0001]$ and $[0,0.1,0.01,0.001]$, respectively. We tune the hyperparameters with $\lambda_{c}\in[0.1,1.0]$, $\lambda_{m}\in[0.1,1.0]$, and $\lambda_{e}\in[0,200]$.
@@ 4.2 Main Results (RQ1)
We compare the performance of our proposed DNR with multiple leading baselines and the results are detailed in Table 1. “DNR-G” and “DNR-B” denote two alternatives to implement $\bm{\epsilon}_{\phi}^{\text{heuristic}}$ by injecting Gaussian noise and Beta noise, respectively, before switching to the model-based sampler. We can observe that the generator-evaluator-based methods (i.e., EGRerank [22], Pier [9], and NAR4Rec [10]) achieve comparable performance among the multiple baselines, validating the effectiveness of the generator-evaluator paradigm in the reranker stage. In comparison, our DNR solutions consistently outperform the state-of-the-art baselines across the three benchmarks. This highlights the effectiveness of our noise reduction formulation and the significance of improving reranker robustness in our solution.
@@ 4.3 Ablation Study (RQ2, RQ3, RQ4, & RQ5)
RQ2: To investigate different ways to leverage retriever scores, we develop three PRM variants for the noise generator: “c score” refers to leveraging a concatenate operation to combine the score features with encoded user states from users’ interaction history; “+ score” denotes incorporating the score feature with user states by a simple addition; “w score” represents adopting the initial retriever scores as the weights of candidate items. The frameworks of the three variants are shown in Appendix C.3. We report the experimental results in Table 2. The variants “c score”, “+ score”, and “w score” achieve better performance than the PRM backbone and slightly better performance than the state-of-the-art baseline Pier, validating that it is beneficial to include the information from retriever scores in the reranking stage. Additionally, our DNR solutions outperform all three variants, indicating a superior strategy to leverage the retriever scores.
Table 2: Ablation Study for the different ways of leveraging the retriever scores. Performance that exceeds Pier’s is indicated with a $\uparrow$ symbol.
@@
RQ3: In Table 1, we also observe that the Beta noise alternative generally performs better in Kuaivideo and Book with significant margin, and the Gaussian noise performs better in ML-1M. This indicates the superiority of Beta over Gaussian is not consistent across recommendation platforms. Nevertheless, we observe that the the learned model-based sampler is more aligned with the Beta distribution in practice (see Appendix D.2).
RQ4: To further investigate different model alternatives, we evaluate the different ways of implementing noise injection and the ablation of learning objectives, and present the results in in Table 3. By comparing “w/ G” with “DNR-G or comparing “w/ B” with “DNR-B”, we can see that the full solution with a well optimized model-based sampler consistently outperforms the heuristic sampler.
RQ5: To compare different learning objectives in Eq.(5), we first observe that “w/ G” and “w/B” methods outperforms the Pier baseline, indicating the effectiveness of $\mathcal{L}_{\text{z}}$, which uses the posterior to augment the learning of reranker. Then, all “w/o $\mathcal{L}_{\text{adv}}$” alternatives and “w/o $\mathcal{L}_{\text{x}}$” alternatives appear to downgrade the effectiveness of the solution, compared to the full implementation of DNR. All these observations validate the effectiveness of the derived terms in Eq.(5).
Table 3: Ablation Study for DNR alternatives. “w/ G” and “w/ B” denote heuristic-only posteriors that adopt noise sampler with Gaussian noise and Beta noise without using model-based sampler; “G w/o $\mathcal{L}_{\text{adv}}$” and “B w/o $\mathcal{L}_{\text{adv}}$” represent DNR-G or DNR-B paradigms without adversarial noise learning objective; “G w/o $\mathcal{L}_{\text{x}}$” and “B w/o $\mathcal{L}_{\text{x}}$” refer to DNR-G or DNR-B paradigms without the noise regularization.

### v2

@@ 3.3 Implementing Noise Generator for the Posterior
  • Heuristic Generator: This option uses a non-parametric noise generator $f_{\phi}^{\text{heuristic}}$ with a manually designed distribution. For example, we can adopt random Gaussian noise similar to Ho et al. (2020); Li et al. (2024), but we notice that this is not a good representation of the noise in retriever scores. Intuitively, given the binary user feedback signals, a natural assumption of the retriever prior is the Beta distribution, where we can derive the heuristic solution as $\bm{\epsilon}\sim\text{Beta}(\alpha,\beta)$. For simplicity, one may set $\alpha=\beta=0.5$ to approximate the uncertainty. In practice, we find that it is also beneficial to set $\alpha$ and $\beta$ that align with the observed noise distribution in the data.
@@
  • Model-based Generator: While the heuristic generator is effective in our empirical study, we are skeptical of whether the heuristic noise generator is optimal for augmenting the retriever score distribution. To achieve a more precise estimation, we design a learnable model $f_{\phi}^{\text{model}}$, with parameter $\phi$, that will be optimized along with the reranker (details in the next section). The generated noise $\bm{\epsilon}$ should also be aligned with the observed distribution, and it also participates in an adversarial training objective in order to better explore the retriever score space, which in turn improves the reranker’s effectiveness. As we will show in Section 4.3, this choice is empirically better than the heuristic generator.
@@ 3.4 Overall Learning Framework
In this section, we show how to optimize the user feedback alignment goal i.e., Eq.(1) by simultaneously learning the denoising reranker and the adversarial noise generator. Specifically, we can decompose the negative log-likelihood objective into three loss terms (derivation in Appendix B.1):
$-\log p(\mathbf{z}_{u})$ $=\mathcal{L}_{\text{z}}+\mathcal{L}_{\text{adv}}+\mathcal{L}_{\text{x}}+\delta_{\text{x}}$ (5)
@@
which consists of an augmented denoising loss term $\mathcal{L}_{\text{z}}$, an adversarial noise generation loss $\mathcal{L}_{\text{adv}}$, and a score distribution regularization term $\mathcal{L}_{\text{x}}$. We will illustrate the details of these three terms in the following sections. The additional term $\delta_{\text{x}}=-D_{\text{KL}}\big(p_{\phi}(\mathbf{x}|\mathbf{z}_{u})\|p_{\text{x|z}}(\mathbf{x}|\mathbf{z}_{u})\big)\leq 0$ represents the negative divergence between the generated noisy retriever scores $p_{\phi}$ and the corresponding real $p_{\text{x|z}}$. This term is not directly optimizable due to the unknown $p_{\text{x|z}}$, but we expect to achieve a small divergence in $\delta_{\text{x}}$ with the optimization of the other three terms, so that the overall optimization is aligned with the goal of Eq.(1).
Denoising with Augmented Retriever Scores: $\mathcal{L}_{\text{z}}$ in Eq.(5) denotes the denoising loss under the noisy retriever scores generated by the noise generator. Specifically, we can first synthesize $\mathbf{x}_{u}^{\prime}\sim p_{\phi}(\cdot|\mathbf{z}_{u})$ as the generated retriever scores according to Eq.(4), then pass them into the reranker $q_{\theta}$ for the prediction of user feedback $\mathbf{z}_{u}$, and $\mathcal{L}_{\text{z}}$ only updates $\theta$. The optimization can also be simplified into the BCE loss with $\mathbf{z}_{u}$ as labels. Note that the generated scores $\mathbf{x}_{u}^{\prime}$, different from the observed $\mathbf{x}_{u}$, essentially serve as the augmented scores that the initial retriever would have produced given the user request. This means that the reranker will be aware of the noisy behavior of the retriever. Combining with the direct optimization, we have the final denoising loss for the reranker:
@@
where $\lambda_{m}$ controls the magnitude of the augmentation, and the parameters of the noise generator $f_{\phi}$ are fixed. Additionally, our model-agnostic framework does not restrict the architecture of the denoising reranker $q_{\theta}$, as we will show in section 4.3, one can adopt various backbones, including PRM Pei et al. (2019) and PIER Shi et al. (2023).
Adversarial Noise Learning: $\mathcal{L}_{\text{adv}}$ describes how the reranker $q_{\theta}$ behaves differently from the user feedback under the viewpoint of the synthetic retriever scores. In practice, this term is expected to be small since the ratio $q_{\theta}(\mathbf{z}_{u}|\mathbf{x})/p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})\approx 1$ in major cases due to the user feedback alignment of Eq.(6) and the augmented objective $\mathcal{L}_{\text{z}}$. Nevertheless, in the view of generated noisy retriever scores, we can see that $\mathcal{L}_{\text{adv}}$ encourages adversarial samples of $\mathbf{x}$ that produce small ratios of $q_{\theta}(\mathbf{z}_{u}|\mathbf{x})/p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})$. In other words, a better $p_{\phi}$ should be able to explore and synthesize retriever scores that are incorrectly denoised by the reranker $q_{\theta}$. For the heuristic noise generator $f_{\phi}^{\text{heuristic}}$, it is not optimizable once $\alpha$ and $\beta$ are chosen. In contrast, for the model-based generator $f_{\phi}^{\text{model}}$, we follow the intuition of adversarial case exploration, and approximate $\mathcal{L}_{\text{adv}}$ by first fixing the reranker and then optimizing the generator with an adversarial loss:
@@
We implement the model-based generator $f_{\phi}^{\text{model}}$ by a 2-layer MLP and directly output $\mathbf{x}_{u}^{\prime}$ using Eq.(4), so that the end-to-end learning becomes feasible for Eq.(7). In cases where the noise generator is not directly optimizable, one may have to consider other alternatives like reparameterization tricks Kingma & Welling (2014) or policy gradient Silver et al. (2014). Notably, $\mathcal{L}_{\text{adv}}$ does not encourage over-exploration of $\mathbf{x}_{u}^{\prime}$, which may negatively impact the learning of the reranker. In such cases, the synthetic scores are unrealistically different from the real retriever scores, diverging from the user feedback, inducing small $p_{\text{z|x}}$ and large $q_{\theta}(\mathbf{z}_{u}|\mathbf{x})/p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})$, violating $\mathcal{L}_{\text{adv}}$.
Noise Regularization: $\mathcal{L}_{\text{x}}$ aims to reduce the divergence between the synthetic retriever score distribution $p_{\phi}$ and the real retriever score prior distribution $p_{\text{x}}$. Intuitively, this term favors the synthetic score distribution $p_{\phi}$ that can mimic the noisy behavior of the real retriever’s scores $p_{\text{x}}$. For implementation, we first synthesize the retriever score $\mathbf{x}_{u}^{\prime}$ with Eq.(4) and calculate the distributional difference between synthetic scores from $p_{\phi}$ and retriever scores from $p_{\text{x}}$. Then, we minimize this difference to approximate the distribution-level KL divergence.
Again, given that the heuristic noise generator $f_{\phi}^{\text{heuristic}}$ is not optimizable by the objectives in Eq (5), it might not be the optimal choice for augmenting the retriever score distribution. As a consequence, this could lead to recommendations that are not in alignment with user feedback. For the training of the model-based noise generator $f_{\phi}^{\text{model}}$, we first adopt the heuristic generator in the first $\lambda_{e}$ epochs to ensure learning stability of reranker $q_{\theta}$ throughout $\mathcal{L}_{z}$, then engage the adversarial learning of $\mathcal{L}_{\text{adv}}$ and the optimization of $\mathcal{L}_{\text{x}}$ afterward, switching the generation of $\bm{\epsilon}$ from heuristic generator to model-based generator. We present the algorithm of our training procedure in Appendix C.
@@ 4 Experiments
RQ1: How does DNR perform in the reranking tasks of recommendation systems compared with leading rerankers? RQ2: How well is the denoising formulation compared to alternative methods of leveraging retriever scores? RQ3: Can the denoising formulation generalize to other reranker backbones, rather than PRM? RQ4: What is the performance difference between the heuristic noise generator and the adversarial model-based generator? RQ5: How does different objectives in DNR contribute to the performance? RQ6: How sensitive is DNR to the hyperparameters $\lambda_{c}$, $\lambda_{m}$, and $\lambda_{e}$?
@@ 4.1 Expermental Settings
Datasets. We conduct experiments on three real-world datasets, including ML-1M Harper & Konstan (2016), Kuaivideo Liu et al. (2025), and Amazon-books Ni et al. (2019). The detailed description and the statistics of the processed datasets for both the retriever and the reranker stages are presented in Appendix D.1. For the online experimental verification, please refer to Appendix E.6.
Baselines: We compare the performance of DNR with four categories of recommendation methods in our reranking task, including traditional recommenders (i.e., SASRec Kang & McAuley (2018), Caser Tang & Wang (2018), GRU4Rec Hidasi et al. (2016), and MiDNN Zhuang et al. (2018)), which individually predict the scores of candidate items and rank them accordingly; list-refinement methods (i.e., DLCM Ai et al. (2018), SetRank Pang et al. (2020), PRM Pei et al. (2019), and MIR Xi et al. (2022)), which encode the mutual inference in candidate items and refine the rankings based on this information; generator-evaluator methods (i.e., EGRerank Huzhang et al. (2023), Pier Shi et al. (2023), and NAR4Rec Ren et al. (2024)), which generate multiple reranked item lists and evaluate them to select the best one for users; and diffusion-Based methods (i.e., DiffuRec Li et al. (2024) and DCDR Lin et al. (2024)) which utilize diffusion models to generate items or lists for recommendation. Detailed information for our baselines is presented in Appendix D.3.
Implementation Details: Our experiments are implemented on Nvidia A40 GPUs with Python 3.8 and PyTorch 1.12. We construct the retriever stage with the collaborative filtering method Koren et al. (2009) to obtain the initial retriever scores for the top-50 candidate items. The detailed setting and experimental results of the retriever stage are presented in Appendix E.1. The reranker reorders these items filtered by retrievers and selects the top $K=6$ for display. The dimension of rerankers’ hidden embedding is 128 across all models, and all models are trained until convergence. The learning rate and weight decay are tuned within the range of $[0.01,0.001,0.0001]$ and $[0,0.1,0.01,0.001]$, respectively. We tune the hyperparameters with $\lambda_{c}\in[0.1,1.0]$, $\lambda_{m}\in[0.1,1.0]$, and $\lambda_{e}\in[0,200]$. The optimal hyperpapemters adopted for our experimental results are presented in Appendix D.2.
@@ 4.2 Main Results (RQ1)
We compare the performance of our proposed DNR with multiple leading baselines and the results are detailed in Table 1. “DNR-G” and “DNR-B” denote using two alternatives of $f_{\phi}^{\text{heuristic}}$ to generate Gaussian noise and Beta noise, respectively, before switching to the model-based noise generator $f_{\phi}^{\text{model}}$ in the epoch $\lambda_{e}$. We can observe that the generator-evaluator-based methods (i.e., EGRerank Huzhang et al. (2023), Pier Shi et al. (2023), and NAR4Rec Ren et al. (2024)) achieve comparable performance among the multiple baselines, validating the effectiveness of the generator-evaluator paradigm in the reranker stage. In comparison, our DNR solutions consistently outperform the state-of-the-art baselines across the three benchmarks. This highlights the superiority of our noise reduction formulation and the significance of improving reranker effectiveness with three objectives in adversarial learning.
@@ 4.3 Ablation Study (RQ2, RQ3, RQ4, & RQ5)
RQ2 & RQ3: To investigate different ways to leverage retriever scores, we develop three variants for leveraging retriever scores as rerankers’ input: “c score” refers to leveraging a concatenate operation to combine the score features with encoded user states from users’ interaction history; “+ score” denotes incorporating the score feature with user states by a simple addition; “w score” represents adopting the initial retriever scores as the weights of candidate items. The frameworks of the three variants are shown in Appendix D.4. To validate the generalizability of DNR across different reranker backbones, we conduct experiments comparing our three variants and DNR itself when using PRM Pei et al. (2019) and Pier Shi et al. (2023) as the underlying rerankers, respectively. We report the experimental results in Table 2. The variants “c score”, “+ score”, and “w score” all achieve superior performance compared to the base reranker backbone, confirming that incorporating retriever scores into the reranking stage yields tangible benefits. Moreover, our DNR outperforms all three variants, underscoring denoising formulation as a more effective strategy for leveraging retriever scores. Importantly, DNR delivers consistent improvements across both backbones, thereby validating the generalizability of our denoising formulation for reranking architectures.
Table 2: Ablation Study for the different ways of leveraging the retriever scores on both PRM and Pier reranker backbones.
@@
DNR 63.41 77.28 68.02 70.22 91.97 48.89 69.35 57.76 56.02 92.38 73.68 82.61 74.82 76.88 93.78
RQ4: To further investigate different noise generators, we conduct ablation experiments on the heuristic noise generators that generate noise from the Gaussian distribution and Beta distribution, denoted as “w/ G” and “w/ B” respectively. The results are presented in Table 3. By comparing “w/ G” with “DNR-G or comparing “w/ B” with “DNR-B”, we can see that our DNR framework—equipped with a model-based noise generator for adversarial learning—consistently outperforms counterparts using heuristic noise generators.
RQ5: To compare different learning objectives in Eq.(5), we first observe that “w/ G” and “w/B” methods outperform the PRM baseline, indicating the effectiveness of $\mathcal{L}_{\text{z}}$, which uses the heuristic noise generator to augment the learning of reranker. Then, all “w/o $\mathcal{L}_{\text{adv}}$” alternatives and “w/o $\mathcal{L}_{\text{x}}$” alternatives appear to downgrade the effectiveness of the full implementation of DNR. All these observations validate the effectiveness of the derived terms in Eq.(5).
Table 3: Ablation Study for DNR alternatives. “G w/o $\mathcal{L}_{\text{adv}}$” and “B w/o $\mathcal{L}_{\text{adv}}$” represent DNR-G or DNR-B paradigms without adversarial noise learning objective; “G w/o $\mathcal{L}_{\text{x}}$” and “B w/o $\mathcal{L}_{\text{x}}$” refer to DNR-G or DNR-B paradigms without the noise regularization.